# Toward Carnot efficient high output power heat engines using bubbly two-phase flow

### v2

@@
Thermodynamic gas power cycles achieving Carnot efficiency require isothermal expansion, which is associated with slow processes and results in negligible power output. This study proposes a practical method for rapid near-isothermal gas expansion, facilitating efficient heat engines without sacrificing power. The method involves bubble expansion in a heat transfer liquid, ensuring efficient and near-isothermal heat exchange. The mixture is accelerated through a converging-diverging nozzle, converting thermal energy into kinetic energy, thereby rotating a reaction turbine for electricity generation. Nozzle experiments with air and water yielded a polytropic index $<$1.052, enabling up to 71% more work extraction than adiabatic expansion. Simulations indicate that utilizing these nozzles for thrust generation enables decreasing heat transfer irreversibilities in the heat engine, consequently resulting in up to 22.6% higher power output than an ideal heat engine based on the organic Rankine cycle. This work paves the way for an efficient and high-power heat-to-power solution.
@@ INTRODUCTION
This study presents the development of an efficient, high-power heat engine based on a novel thermodynamic vapor cycle incorporating quasi-isothermal expansion of bubbles in a heat transfer liquid (HTL). The large bubbles’ surface area supports efficient heat transfer between the phases, and the substantial mass and heat capacity discrepancies between the phases facilitate minimal temperature drop, thereby implementing a quasi-isothermal process. The two-phase mixture flows through converging-diverging nozzles, converting the thermal energy to thrust, which rotates a reaction turbine. To this end, we combine experimental pressure measurements and computational fluid dynamic (CFD) analysis of a two-phase flow in a converging-diverging nozzle to extract an upper boundary for the polytropic index of air and quantify the work enhancement of the near-isothermal compared to adiabatic expansion. Building on this, we simulate the performance of a heat engine that integrates the nozzles as expanders to drive a reaction turbine for electricity generation from a constant-temperature heat source. The results highlight the potential of this design, as a polytropic index of air $n<$1.052 is determined, resulting in work extraction $>$91% of perfect isothermal work. Thermodynamic simulations indicate that the proposed heat engine produces up to 22.6% greater power output than an ORC-based counterpart for a 100–374 ${}^{\circ}C$ constant temperature heat source. Importantly, the isothermal expansion of the cycle facilitates the use of ’wet’ fluids such as water—fluids that challenge the performance in ORC systems but are well-matched for high-temperature heat sources. Since the turbine’s practical efficiency is obscure, we demonstrate that turbine efficiencies between 70% and 90% yield an improvement of 42.7% and 90.7%, respectively, compared to the ORC heat engines with turbine isentropic efficiency of 75% at the temperature of 373 ${}^{\circ}C$. These findings provide a basis for developing efficient heat engines tailored to heat-to-power applications.
@@
The nozzle is designed according to a previously developed model for two-phase flow in a converging-diverging nozzle 21. The design is based on the homogenous model of a bubbly mixture flowing in the nozzle, which was shown valid by others 22, 23, 24, 17, 18. Under this model, we assume that the phases are in pressure, temperature, and velocity equilibrium. The geometry of the cross-section is selected as circular. In contrast to the propulsion nozzle, where the pressure at the injection chamber is dictated and limited by the cruise velocity, the inlet pressure is a free parameter controlled by a pump, which results in supersonic flow. To illustrate the physical principle and evaluate the expansion, we introduce air into a water stream at ambient temperature. The injection pressure at the design point was chosen to be 6 bar, an air mass flow rate of 13.1 $g/s$, an injection temperature of 26 ${}^{\circ}C$, and a water volumetric flow rate of 14.6 $m^{3}/hr$. The nozzle is designed to implement a constant $dp/dt$ to ease the heat transfer between the phases. Figure 1 shows a schematic representation of the designed nozzle.
@@
Subsequently, we employ the polytropic index to gauge the extent to which this process approximates a strictly isothermal expansion. To quantify the polytropic index, we apply the energy conservation equation to the nozzle’s outlet, where the pressure is P, and the mixture’s velocity is expressed by:
$u^{2}=\frac{2\dot{V}_{w}(P_{inj}-P)}{\dot{m}_{w}+\dot{m}_{a}}+u_{inj}^{2}+\frac{2}{\dot{m}_{w}+\dot{m}_{a}}\times(\dot{W}_{a}-\dot{E}_{loss})$ (1)
where $\dot{V}_{w},P_{inj},\dot{m}_{a},\dot{m}_{w},u_{inj},\dot{W}_{a}$ and $\dot{E}_{loss}$ are the water volumetric flow rate, air injection pressure, air mass flow rate, water mass flow rate, water injection velocity, work rate exerted by air, and energy losses, respectively. Further discussion on the governing equations can be found in the experimental procedure section. The air expansion work is assumed to be a polytropic process, for which the work rate is calculated by 12:
@@
where n and $P_{out}$ are the polytropic index and the outlet pressure, respectively. As a result, equation 1 involves two unknown variables: the polytropic index and the head losses in the nozzle.
To gain insights on the head losses in the nozzle, we evaluate the expression $\dot{E}_{loss}=\dot{V}_{w}\Delta P$, where $\Delta P$ is the pressure loss, and is calculated by $\Delta P=\zeta\frac{\rho u^{2}}{2}$. Here, $\zeta$, $\rho$, and $u$ are the resistance coefficient, mixture’s density and inlet velocity, respectively 27.
@@
Figure 5 shows the isothermal efficiency using the extracted polytropic index of $n=1.052$ for pressure ratios up to 10, for which isothermal efficiencies $>94\%$ are attained. As such, an additional 44-71 % more work can be extracted compared to adiabatic expansion, for which $n=1.4$. In the context of heat engines using organic working fluids, these are associated with a lower polytropic index of $n=1.05-1.3$. As such, increasing work extraction by approaching isothermal expansion is limited and dependent on the polytropic index of the working fluid.
@@
The near-isothermal expansion of bubbles experienced in our nozzle reveals opportunities for a new type of heat engine utilizing such nozzles for a reaction turbine. The turbine consists of an impeller connected to nozzles. The rotor leg is submerged in the HTL reservoir, sucking HTL into the rotor. Similarly to a centrifugal pump impeller, the impeller vanes pressurize the HTL until reaching maximal static pressure at the end of the impeller, where the nozzles begin, and liquid WF is injected and mixed with the HTL. However, unlike the centrifugal pump, our impeller has no stator, which minimizes energy losses while maintaining high design flexibility. In the nozzles, heat is transferred from the HTL to the WF until the WF is fully evaporated, and the mixture is accelerated until it reaches the nozzles’ outlet, where the jet flows tangentially, creating a thrust force and rotates the shaft for electricity generation. An example of the impeller design is shown in Figure 6.
@@
The heat source passes through a heat exchanger (1$\rightarrow$2), which transfers heat to the HTL (3$\rightarrow$4). The HTL is pumped by the turbine in conjunction with colder HTL from the reservoir (5). The HTL flow from the reservoir is used to obtain the desired HTL flow rate in the nozzle, resulting in a specific void fraction. Before reaching the nozzle, the impeller increases the HTL’s pressure to the WF’s saturation pressure. The HTL reaches the nozzle’s inlet and mixes with colder WF liquid (6) in the nozzle’s injection chamber. Heat transfers to the WF, which raises its temperature to its evaporation temperature. The WF then completely evaporates and expands inside the nozzle within the HTL until it reaches its outlet (7). This expansion raises the mixture’s velocity and generates thrust, which turns the reaction turbine. The rotor rotates in a tank that functions as a vertical vapor-liquid separator, separating the WF from the HTL after the expansion by generating a film flow of the mixture on the tank boundaries. The hot WF gas enters a recuperator, which transfers heat from the hot WF to the cold condensed WF before it enters the condenser (8$\rightarrow$9). This internal regeneration, facilitated by the isothermal expansion, is used to heat the cold liquid WF, minimizing heat transfer irreversibilities. Then, the liquid WF is pumped through the recuperator (10$\rightarrow$11). The WF is then injected into the nozzles, mixing with the hot HTL (6).
@@
Therefore, the ISO cycle was simulated with water as the WF and a heat source in 100-374 C, covering a substantial portion of the waste heat recovery temperature range. The simulated heat source is characterized by a constant temperature, representing condensation scenarios harnessing constant-temperature latent heat. This application is optimally suited for the ISO cycle, as the predominant portion of the heat involved is latent heat, allowing significant heat extraction with minimal temperature change. The scale of the system is a few hundred kW, which is the scale needed for many industrial waste heat recovery systems. The heat sink temperature was chosen as 20 C, typical for atmospheric conditions.
@@
Since the reaction turbine has yet to be built and experimented with, we consider turbine efficiencies in the range of 20-90%, where the ISO turbine efficiency is defined by the ratio of the turbine output power to the WF expansion work. In this context, we assume the entropy produced due to turbine irreversibilities is retained within the system in heat. In contrast to the ORC turbine, where the generated entropy cannot be utilized, the ISO cycle incorporates this thermal energy into the HTL stream, thereby augmenting the flow rate of the working fluid. This process mitigates the decline in work output due to turbine inefficiencies. Figure 11 illustrates the output power enhancement of the ISO cycle, using water as the working fluid, across the varying turbine efficiencies relative to the ORC, considering a realistic ORC expander exhibits isentropic efficiency of 75%, consistent with values used in past studies 4. The black curve in the figure denotes the ISO turbine efficiency for which the output power of the two cycles is equal. The figure indicates that in the considered ISO turbine efficiencies, the ISO cycle potentially generates up to 90.7% more power than the ORC-based heat engine and is superior to the ORC-based heat engines for ISO turbine efficiencies exceeding 47% in the examined temperature range.
@@ Conclusions
This study introduces the concept of a supersonic two-phase nozzle as a key component of a novel thermodynamic cycle utilizing isothermal expansion. Experimental pressure measurements with air in water validated the supersonic conditions, while CFD allowed us to extract an upper bound for the polytropic index of air of $n<1.052$. This index indicates quasi-isothermal expansion, enabling up to 71% more work extraction than adiabatic expansion. Compared to other reported methods, such as flooded expansion, this method achieves a much higher liquid-to-working-fluid mass ratio (0.997 compared to 0.3). This significantly increases the heat transfer between the phases, enabling almost isothermal gas expansion. Furthermore, using the two-phase nozzle is an important step for increasing the pressure ratio, which makes it more suitable for various applications.
To leverage this nozzle for realizing a heat engine, we theoretically explored a reaction turbine, which rotates due to thrust created by the nozzles. CFD analysis of the impeller shows hydraulic efficiency of $\sim 97\%$, supporting its efficiency potential compared to existing small-scale turbines.
@@
We broadened our simulation to encompass turbine efficiencies spanning from 20% to 90% and compared to ORC small-scale heat engine, which encompasses turbine isentropic efficiency of 75%. This analysis revealed that the ISO cycle outperforms the ORC within the examined temperature range when turbine efficiency exceeds 47%.
Not only is it more efficient, but the ISO turbine has the potential to be much smaller than an ORC counterpart. This is because the energy carrier in the ISO cycle is liquid, three orders of magnitude denser than gas, which is the typical energy carrier in ORC and other heat engines.
@@ ACKNOWLEDGMENTS
All authors are employees and shareholders of Lava Energy Ltd. Carmel Rotschild is a founder of Lava Energy and a member of its scientific advisory board. The authors have submitted the following relevant patents:

### v3

@@
Thermodynamic gas power cycles achieving Carnot efficiency require isothermal expansion, which is associated with slow processes and results in negligible power output. This study proposes a practical method for rapid near-isothermal gas expansion, facilitating efficient heat engines without sacrificing power. The method involves bubble expansion within a heat transfer liquid, ensuring efficient and near-isothermal heat exchange. The mixture is accelerated through a converging-diverging nozzle, converting thermal energy into kinetic energy, thereby rotating a reaction turbine for electricity generation. Experiments with air and water yielded a polytropic index $<$1.052, enabling up to 71% more work extraction than adiabatic expansion. Simulations indicate that utilizing these nozzles for thrust generation decreases heat transfer irreversibilities in the heat engine, resulting in up to 22.6% higher power output than an ideal heat engine based on the organic Rankine cycle. This work paves the way for an efficient and high-power heat-to-power solution.
@@ INTRODUCTION
This study presents the development of an efficient, high-power heat engine based on a novel thermodynamic vapor cycle incorporating quasi-isothermal expansion of bubbles in a heat transfer liquid (HTL). The large bubbles’ surface area supports efficient heat transfer between the phases, and the substantial mass and heat capacity discrepancies between the phases facilitate minimal temperature drop, thereby implementing a quasi-isothermal process. The two-phase mixture flows through converging-diverging nozzles, converting the thermal energy to thrust, which rotates a reaction turbine. To this end, we combine experimental pressure measurements and computational fluid dynamic (CFD) analysis of a two-phase flow in a converging-diverging nozzle to extract an upper boundary for the polytropic index of air and quantify the work enhancement of the near-isothermal compared to adiabatic expansion. Building on this, we simulate the performance of a heat engine that integrates the nozzles as expanders to drive a reaction turbine for electricity generation from a constant-temperature heat source. The results highlight the potential of this design, as a polytropic index of air $n<$1.052 is determined, resulting in work extraction $>$91% of perfect isothermal work. Thermodynamic simulations indicate that the proposed heat engine produces up to 22.6% greater power output than an ORC-based counterpart for a 100–374 ${}^{\circ}C$ constant temperature heat source. Importantly, the isothermal expansion of the cycle facilitates the use of ’wet’ fluids such as water—fluids that challenge the performance in ORC systems but are well-suited for high-temperature heat sources. Since the turbine’s practical efficiency is obscure, we demonstrate that turbine efficiencies between 70% and 90% yield an improvement of 42.7% and 90.7%, respectively, compared to the ORC heat engines with turbine isentropic efficiency of 75% at the temperature of 373 ${}^{\circ}C$. These findings provide a basis for developing efficient heat engines tailored to heat-to-power applications.
@@
The nozzle is designed according to a previously developed model for two-phase flow in a converging-diverging nozzle 21. The design is based on the homogenous model of a bubbly mixture flowing in the nozzle, which was shown valid by others 22, 23, 24, 17, 18. Under this model, we assume that the phases are in pressure, temperature, and velocity equilibrium. In contrast to the nozzles designed for underwater propulsion where the pressure at the injection chamber is dictated and limited by the cruise velocity, the inlet pressure is a free parameter controlled by a pump, which results in supersonic flow.
To illustrate the physical principle and evaluate the expansion, we introduce air into a water stream at ambient temperature. The injection pressure at the design point was chosen to be 6 bar, an air mass flow rate of 13.1 $g/s$, an injection temperature of 26 ${}^{\circ}C$, and a water volumetric flow rate of 14.6 $m^{3}/hr$. The geometry of the cross-section is selected as circular. The nozzle is designed to implement a constant $dp/dt$ to ease the heat transfer between the phases. Figure 1 shows a schematic representation of the designed nozzle.
@@
Subsequently, we employ the polytropic index to gauge how much this process approximates a strictly isothermal expansion. To quantify the polytropic index, we apply the energy conservation equation to the nozzle’s outlet, where the pressure is P, and the mixture’s velocity is expressed by:
$u_{out}^{2}=\frac{2\dot{V}_{w}(P_{inj}-P_{out})}{\dot{m}_{w}+\dot{m}_{a}}+u_{inj}^{2}+\frac{2}{\dot{m}_{w}+\dot{m}_{a}}\times(\dot{W}_{a}-\dot{E}_{loss})$ (1)
where $\dot{V}_{w},P_{inj},P_{out},\dot{m}_{a},\dot{m}_{w},u_{inj},u_{out},\dot{W}_{a}$ and $\dot{E}_{loss}$ are the water volumetric flow rate, air injection pressure, outlet pressure, air mass flow rate, water mass flow rate, mixture’s velocity at injection, mixture’s velocity at nozzle’s outlet, work rate exerted by air, and energy losses, respectively. Further discussion on the governing equations can be found in the experimental procedure section. The air expansion work is assumed to be a polytropic process, for which the work rate is calculated by 12:
@@
where n is the polytropic index. As a result, equation 1 involves two unknown variables: the polytropic index and the head losses in the nozzle.
To gain insights on the energy losses in the nozzle, we evaluate the expression $\dot{E}_{loss}=\dot{V}_{w}\Delta P$, where $\Delta P$ is the head loss, and is calculated by $\Delta P=\zeta\frac{\rho u^{2}}{2}$. Here, $\zeta$, $\rho$, and $u$ are the resistance coefficient, mixture’s density and inlet velocity, respectively 27.
@@
Figure 5 shows the isothermal efficiency using the extracted polytropic index of $n=1.052$ for pressure ratios up to 10, for which isothermal efficiencies $>94\%$ are attained. As such, an additional 44-71 % more work can be extracted compared to adiabatic expansion, for which $n=1.4$.
@@
The near-isothermal expansion of bubbles experienced in our nozzle reveals opportunities for a new type of heat engine utilizing such nozzles for a reaction turbine. The turbine consists of an impeller connected to nozzles. The rotor leg is submerged in the HTL reservoir, sucking HTL into the rotor. Similarly to a centrifugal pump impeller, the impeller vanes pressurize the HTL until reaching maximal static pressure at the end of the impeller, where the nozzles begin, and liquid WF is injected and mixed with the HTL. However, unlike the centrifugal pump, our impeller has no stator, minimizing energy losses while maintaining high design flexibility. In the nozzles, heat is transferred from the HTL to the WF until the WF is fully evaporated, and the mixture is accelerated until it reaches the nozzles’ outlet, where the jet flows tangentially, creating a thrust force and rotates the shaft for electricity generation. An example of the impeller design is shown in Figure 6.
@@
The heat source flows through a heat exchanger (1$\rightarrow$2), transferring heat to the HTL (3$\rightarrow$4). The HTL is pumped by the turbine in conjunction with colder HTL from the reservoir (5). The HTL flow from the reservoir is used to obtain the desired HTL flow rate in the nozzle, resulting in a specific void fraction. Before reaching the nozzle, the impeller increases the HTL’s pressure to the WF’s saturation pressure. The HTL reaches the nozzle’s inlet and mixes with colder WF liquid (6) in the nozzle’s injection chamber. Heat transfers to the WF, which raises its temperature to its evaporation temperature. The WF then completely evaporates and expands inside the nozzle within the HTL until it reaches its outlet (7). This expansion raises the mixture’s velocity and generates thrust, which turns the reaction turbine. The rotor rotates in a tank that functions as a vertical vapor-liquid separator, separating the WF from the HTL after the expansion by generating a film flow of the mixture on the tank boundaries. The hot WF gas enters a recuperator, transferring heat from the hot WF to the cold condensed WF before it enters the condenser (8$\rightarrow$9). This internal regeneration, facilitated by the isothermal expansion, is used to heat the cold liquid WF, minimizing heat transfer irreversibilities. Then, the liquid WF is pumped through the recuperator (10$\rightarrow$11). The WF is then injected into the nozzles, mixing with the hot HTL (6).
@@
Therefore, the ISO cycle was simulated with water as the WF and a heat source in 100-374 C, covering a substantial portion of the waste heat recovery temperature range. The simulated heat source is characterized by a constant temperature, representing condensation scenarios harnessing constant-temperature latent heat. This application is optimally suited for the ISO cycle, as the predominant portion of the heat involved is latent heat, allowing significant heat extraction with minimal temperature change. The scale of the system is a few hundred kW, which is the scale needed for many industrial waste heat recovery systems. The heat sink temperature was chosen as 20 C.
@@
Since the reaction turbine has yet to be built and experimented with, we consider turbine efficiencies in the range of 20-90%, where the ISO turbine efficiency is defined by the ratio of the turbine output power to the WF expansion work. In this context, we assume the heat produced due to turbine irreversibilities is retained within the system. In contrast to the ORC turbine, where the generated heat cannot be utilized, the ISO cycle incorporates this thermal energy into the HTL stream, thereby augmenting the flow rate of the working fluid. This process mitigates the decline in work output due to turbine inefficiencies. Figure 11 illustrates the output power enhancement of the ISO cycle using water as the working fluid across the varying turbine efficiencies relative to the ORC. In this analysis, we consider a realistic ORC expander isentropic efficiency of 75%, consistent with values used in past studies 4. The black curve in the figure denotes the ISO turbine efficiency for which the output power of the two cycles is equal. The figure indicates that at the considered ISO turbine efficiencies range, the ISO cycle potentially generates up to 41% more power than the ORC-based heat engine and is superior to the ORC-based heat engines for ISO turbine efficiencies exceeding 65% in the examined temperature range.
@@ Conclusions
This study introduces the concept of a supersonic two-phase nozzle as a key component of a novel thermodynamic cycle utilizing isothermal expansion. Experimental pressure measurements with air in water validated the supersonic conditions, while CFD allowed us to extract an upper bound for the polytropic index of air of $n<1.052$. This index indicates near-isothermal expansion, enabling up to 71% more work extraction than adiabatic expansion. Compared to other reported methods, such as flooded expansion, this method achieves a much higher liquid-to-working-fluid mass ratio (0.997 compared to 0.3). This significantly increases the heat transfer between the phases, enabling almost isothermal gas expansion. Furthermore, using the two-phase nozzle is an important step for increasing the pressure ratio, making it more suitable for various applications.
To leverage this nozzle for realizing a heat engine, we theoretically explored a reaction turbine, rotating due to thrust created by the nozzles. CFD analysis of the impeller shows hydraulic efficiency of $\sim 97\%$, supporting its efficiency potential compared to existing small-scale turbines.
@@
We broadened our simulation to encompass turbine efficiencies spanning from 40% to 90% and compared to ORC small-scale heat engine, which encompasses turbine isentropic efficiency of 75%. This analysis revealed that the ISO cycle outperforms the ORC within the examined temperature range when turbine efficiency exceeds 65%.
Not only is it more efficient, but the ISO turbine also has the potential to be significantly smaller than its ORC counterpart. This is because the energy carrier in the ISO cycle is liquid, three orders of magnitude denser than gas, which is the typical energy carrier in ORC and other heat engines.
@@ ACKNOWLEDGMENTS
All authors are employees and stakeholders of Lava Energy Ltd. Carmel Rotschild is a founder of Lava Energy and a member of its scientific advisory board. The authors have submitted the following relevant patents: